OpenStudy (anonymous):

Find the derivative of cos(arcsin 2x)?

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left\{ \cos \left( \arcsin2x \right) \right\}=-\sin \left( arc \sin 2x \right)*\frac{ 1 }{ \sqrt{1-4x^2} }*2\]

OpenStudy (mathmale):

@Surjithayer: with all due respect, i feel I must ask you again to please NOT provide answers. Our goal here is to help the learner find his or her own solutions. Indeed you know this material, and you know it well; please turn your skills towards guiding others as they do their own work. Thank you. MM